# Training-Free Text-Guided Color Editing with Multi-Modal Diffusion Transformer

### v2

@@ B.3. Limitations
Figure 12. Examples of failure cases. Left: results generated with SD3. Right: results generated with FLUX.1-dev.
@@
Figure 13. Qualitative image results compared with training-free methods and commercial models on ColorCtrl-Bench. The left four columns are generated using FLUX.1-dev, while the right three are generated using SD3. Best viewed with zoom-in.
Figure 14. Qualitative video results compared with training-free models on ColorCtrl-Bench. Left: “yellow bus” $\to$ “green yellow”. Right: “green backpack” $\to$ “yellow backpack”. Each shows three frames.
Figure 15. Additional qualitative image results compared with training-free methods and commercial models on PIE-Bench. The left three columns are generated using FLUX.1-dev, while the right two are generated using SD3. Best viewed with zoom-in.
Figure 16. Additional qualitative video results compared with training-free models on PIE-Bench. Left: “white shirt” $\to$ “blue shirt”. Right: “white lamb” $\to$ “blue lamb”. Each shows three frames.
Figure 17. The prompt used to generate ColorCtrl-Bench with GPT.

### v3

@@ B.10 Limitations
Figure 16: Examples of failure cases. Left: results generated with SD3. Right: results generated with FLUX.1-dev.
@@
Appendix C Large Language Model (LLM) Usage
We used an LLM to correct grammar, to draft the web UI for the user study, and to generate the ColorCtrl-Bench prompt list. All outputs were manually reviewed and edited by the authors.
Figure 17: Qualitative image results compared with training-free methods and commercial models on ColorCtrl-Bench. The left four columns are generated using FLUX.1-dev, while the right three are generated using SD3. Best viewed with zoom-in.
Figure 18: Additional qualitative image results compared with training-free methods and commercial models on PIE-Bench. The left three columns are generated using FLUX.1-dev, while the right two are generated using SD3. Best viewed with zoom-in.
Figure 19: Qualitative video results compared with training-free models on ColorCtrl-Bench. Left: “yellow bus” $\to$ “green yellow”. Right: “green backpack” $\to$ “yellow backpack”. Each shows three frames.
Figure 20: Additional qualitative video results compared with training-free models on ColorCtrl-Bench. Left: “gray sofa” $\to$ “brown sofa”. Right: “brown suitcase” $\to$ “black suitcase”. Each shows three frames.
Figure 21: Additional qualitative video results compared with training-free models on PIE-Bench. Left: “white shirt” $\to$ “blue shirt”. Right: “white lamb” $\to$ “blue lamb”. Each shows three frames.
Figure 22: User interface for user study.
Figure 23: The prompt used to generate ColorCtrl-Bench with GPT.